1. Do you see something conceptually wrong with this table?  In other words, the logic of accelerated depreciation is that we derive greater benefits from the asset in early years.  According to that logic, what is wrong with this schedule?

2. Can you propose a solution or modification to the DDB depreciation schedule that would solve the problem you identified in question 1, but would still largely follow the DDB approach?year 1 year 2 DDB over 5 years (not tax) 40.00% 24.00% 14.40% 8.64% 12.96% 100.00% End of year basis 60.00% 36.00% 21.60% 12.
Answer #1

1)In case of double declining depreciation method ,depreciation rate = 2/ useful life

                                                                                                                       2/ 5 = .40 or 40%

So for every year ,depreciation expense = cost (or carrying value) *depreciation rate.

Yes,there is a mistake in table provided .In year 5 ,depreciation charged is equals to 100% of carrying value which is wrong .However depreciation expense should be 12.96 *40%

                                                      = 5.184 %

leaving end of year basis = 12.96-5.184 = 7.776%

2)Modified table will looks like:

DDB over 5 years End of year basis
1 40% 60%
2 24% 36%
3 14.40% 21.60%
4 8.64% 12.96%
5 12.96* 40%= 5.184% 7.776%
Total 92.224%